# A Bayesian latent Gaussian conditional autoregressive copula model for analyzing spatially-varying trends in rainfall

### v1

@@ 5.2 MCMC implementation
We implement the MCMC algorithm in R (http://www.r-project.org). For both simulation studies and the Indian rainfall data application, we draw 200,000 samples from the posterior distribution and discard the first 40,000 observations as burn-in. Further, we perform thinning by 20 iterations and draw inferences based on 8,000 post-burn-in samples from the posterior distribution. For the real data application, where the data includes missing observations at 3 out of 64 years, the computation time is approximately 36 minutes. In the simulation studies, which do not contain any missing data, the computation time is approximately 18 minutes. We perform all the computations on a workstation with an AMD Ryzen 9 5900x processor and 64GB RAM. We monitor the convergence and mixing through Geweke statistics and effective sample sizes [59].
@@ 6 Simulation study
We perform a simulation study focusing on the correctness of the estimation procedure detailed in Section 5 and also to study how spatial correlation can effectively help in reducing the uncertainty of the parameter estimates. For each of the three choices of $\rho\in\{0,0.5,0.9\}$, representing zero, moderate, and high spatial correlation, we simulate 100 datasets from the proposed model in Section 4.2. Here, we set the true parameter values for $\bm{a}$, $\bm{b}$, and $\bm{c}$ to the estimates obtained from the real dataset in Figure 2. Here, we compare models having a CAR structure at the data level and either independence, CAR, or ICAR structure at the prior level. For all values of $\rho$ in the data-generating models, the marginal distribution of the observations remains the same for each region due to choosing the same areally-varying coefficient vectors.
@@
For each model, we compare the Mean Square Errors (MSEs), posterior Standard Deviations (SDs), and Coverage Probabilities (CovPs) of the parameters $\bm{a},\bm{b},\bm{c}$ and $\rho$ based on the posterior samples drawn using MCMC. For each of the 100 datasets, we obtain MSEs and posterior SDs for each $a_{i}$, $b_{i}$, and $c_{i}$. Here, we report the average of the MSEs for all 34 $a_{i}$s and 100 datasets, for example. For $\rho$, we report the average of the MSEs for all 100 datasets. We also obtain the SDs reported here via a similar averaging. For calculating CovPs, we first calculate the 95% credible interval for the corresponding parameters and check whether the true value falls within that interval. Thus, we obtain a binary value for each dataset and each $i$. Further, we average across all 34 regions and 100 datasets for $a_{i}$s, $b_{i}$s, and $c_{i}$s and report the average. For $\rho$, we only obtain a similar binary value from each dataset and report the average across 100 datasets. We report the results in Table 1. When the true data generating model has $\rho=0$, we leave the CovPs blank for $\rho$ because the posterior distribution is supported over $(0,1)$ and does not include zero.
@@
When true $\rho=0$, comparing all the metrics across models for $a_{i}$s, we observe that the CAR-ICAR model provides the smallest MSE, the smallest SD, and the highest CovPs. We observe similar results for $b_{i}$s, $c_{i}$s, and $\rho$, indicating that an ICAR prior is more suitable due to the natural spatial pattern of the true areally-varying coefficients obtained from the Indian rainfall dataset. Here, the posterior variances of the parameters $\rho_{a}$, $\rho_{b}$, and $\rho_{c}$ are high. Hence, even if an ICAR prior is less flexible (due to fixed spatial autocorrelation one), it provides a better model fitting. When true $\rho=0.5$, i.e., when the spatial correlation at the data layer is moderate, we see a similar performance comparison for $a_{i}$s and $b_{i}$s. For $c_{i}$s, while the average MSE is the smallest for the CAR-ICAR model, the CAR-Indep model provides the smallest average posterior SD, although comparable with the CAR-ICAR model. For $\rho$, in terms of MSE, CAR-Indep and CAR-CAR models perform the best jointly, while for the CAR-Indep model, the average posterior SD is the smallest. When true $\rho=0.9$, i.e., when the spatial correlation at the data layer is strong, the CAR-ICAR and CAR-CAR models usually provide better results for all metrics, except for the low coverage probability (0.9) for $\rho$ in the case of CAR-ICAR and CAR-CAR models. Based on a global comparison with all data-generating scenarios, the CAR-ICAR model is parsimonious and performs equally or better than the alternatives.
@@
Table 1: Average MSE, posterior SD, and coverage probability (CovP) of the areally-varying coefficients and the spatial autocorrelation parameter for independent, CAR, and ICAR priors along with a CAR model specification for the data layer. Here, we consider low through high values of the true spatial autocorrelation parameter and report the results averaged across different regions and 100 simulated datasets (for areally-varying coefficients) and for $\rho$ by averaging across 100 simulated datasets. A smaller MSE and SD value and a higher CovP value are preferred.
CAR-Indep CAR-ICAR CAR-CAR
@@
$\rho=0$ $\bm{a}$ 48.283 4.405 0.946 43.449 4.164 0.952 44.7 4.244 0.951
$\bm{b}$ 1.665 2.791 0.95 1.653 2.784 0.947 1.66 2.787 0.949
$\bm{c}$ 1.162 2.31 0.95 1.073 2.203 0.946 1.115 2.278 0.95
$\rho$ 5.327 4.06 - 5.105 4.005 - 5.185 4.014 -
$\rho=0.5$ $\bm{a}$ 46.165 4.32 0.941 41.638 4.106 0.945 42.737 4.176 0.949
$\bm{b}$ 1.716 2.789 0.945 1.709 2.785 0.947 1.712 2.788 0.946
$\bm{c}$ 1.216 2.272 0.931 1.153 2.273 0.943 1.175 2.295 0.945
$\rho$ 4.109 4.474 0.94 4.111 4.485 0.96 4.109 4.487 0.96
$\rho=0.9$ $\bm{a}$ 40.089 3.991 0.944 37.304 3.867 0.947 37.966 3.904 0.948
$\bm{b}$ 1.799 2.783 0.937 1.768 2.779 0.94 1.776 2.78 0.941
$\bm{c}$ 1.185 2.24 0.93 1.203 2.403 0.959 1.17 2.339 0.95
$\rho$ 0.55 1.804 0.94 0.552 1.797 0.9 0.552 1.796 0.9
[…]
@@ 7 Indian rainfall data analysis
We first fit the CAR-Indep, CAR-CAR, and CAR-ICAR models to the Indian rainfall dataset and compare the models based on specific metrics. Here, unlike the simulation settings, the true values of the parameters are not known, and hence, we cannot compare the models in terms of MSEs and coverage probabilities. Hence, we only compare the average posterior SDs for $a_{i}$s, $b_{i}$s, and $c_{i}$s and the posterior SD of $\rho$. Besides, we compare the DIC and WAIC of the models as described in Section 2.6. We report the results in Table 2. Except for the average posterior SD of $c_{i}$s, all other metrics indicate a superior performance of the CAR-ICAR model. Regarding DIC and WAIC, the ICAR prior provides a better fitting and prediction performance than the CAR prior, followed by the independent priors; this result indicates the suitability of allowing spatially-varying coefficients in a Gaussian graphical model framework.
@@
$\frac{1}{n}\sum_{i}sd(a_{i})$ $\frac{1}{n}\sum_{i}sd(b_{i})$ $\frac{1}{n}\sum_{i}sd(c_{i})$ $sd(\rho)$ DIC WAIC
Ind 4.105 2.753 2.077 1.225 28149.7 2235948
ICAR 3.887 2.786 2.293 1.141 28098.7 2201614
CAR 3.983 2.844 2.21 1.198 28128.9 2240857
[…]
@@
We further consider the CAR-ICAR model only for drawing inferences and report results obtained based on it. To assess the quality of MCMC chains, we use Geweke statistics for the convergence of the post-burn-in samples and the effective number of samples (ESS) for mixing and sample autocorrelation. We report the average of the Geweke statistics and the ESS values for areally-varying coefficients. A high value of ESS is preferred. The Geweke statistic is a $z$-score; hence, its absolute value being lower than 1.96 indicates the convergence of the MCMC chains. We report the results in Table 3. While the ESS values are relatively higher for all the model hyperparameters, they are usually smaller for the main model parameters $\bm{a}$, $\bm{b}$, $\bm{c}$, and $\rho$. Overall, the ESS values are reasonable for all parameters and hyperparameters. Further, the Gewke statistics for all parameters and hyperparameters are close to zero and not outside the interval ±$\pm 1.96$, indicating convergence of the MCMC chains after the burn-in periods. Altogether, both the metrics suggest that the MCMC samples are of reasonably good quality for drawing statistical inferences.
@@
Table 3: Average effective sample size and Geweke statistics of different parameters of the CAR-ICAR model
@@
We report the posterior means and SDs of $a_{i}$s, $b_{i}$s, and $c_{i}$s in Figure 3. Besides, the posterior mean and SD of $\rho$ are 0.9309 and 0.0114, respectively. The spatial patterns of the coefficients are similar to the ones observed in Figure 2, which indicates that the CAR-ICAR model can capture the spatial patterns of the coefficients correctly. The posterior means of $a_{i}$s, i.e., the shape parameters of the gamma model, are higher in the middle, eastern, and southern parts of India. At the same time, they are low in the northwestern parts of India and also near the northeastern meteorological subdivisions. Some eastern and western Indian subdivisions showcase higher posterior SD than others. The posterior means of $b_{i}$s, the intercept-related parameters of the gamma model, are higher in the northwestern and northeastern meteorological subdivisions. At the same time, they are lowest in the middle parts of the country. Finally, a positive $c_{i}$ indicates a negative trend follows from (10). We observe positive estimates of $c_{i}$s throughout the country, indicating a negative trend in the amount of annual total rainfall. However, due to high posterior SDs, most of these estimates are insignificant except for three meteorological subdivisions: 1. Arunachal Pradesh, 2. Assam and Meghalaya, and 3. Nagaland, Manipur, Mizoram, and Tripura. A significantly negative trend for these Himalayan regions indicates a mean annual total rainfall decline and thus requires further environmental assessment.

### v2

@@ 5.2 MCMC implementation
We implement the MCMC algorithm in R (http://www.r-project.org). For both simulation studies and the Indian rainfall data application, we draw 200,000 samples from the posterior distribution and discard the first 40,000 observations as burn-in. Further, we perform thinning by 20 iterations and draw inferences based on 8,000 post-burn-in samples from the posterior distribution. For the real data application, where the data includes missing observations at 3 out of 64 years, the computation time is approximately 36 minutes. In the simulation studies, which do not contain any missing data, the computation time is approximately 18 minutes. We perform all computations on a workstation equipped with an AMD Ryzen 9 5900X processor and 64GB of RAM. We monitor the convergence and mixing through Geweke statistics and effective sample sizes [reich2019bayesian].
@@ 6 Simulation study
We conduct a simulation study to assess the accuracy of the estimation procedure outlined in Section 5 and to investigate how spatial correlation can effectively reduce the uncertainty of parameter estimates. For each of the three choices of $\rho\in\{0,0.5,0.9\}$, representing zero, moderate, and high spatial correlation, we simulate 100 datasets from the proposed model in Section 4.2. Here, we set the true parameter values for $\bm{a}$, $\bm{b}$, and $\bm{c}$ to the estimates obtained from the real dataset in Figure 2. Here, we compare models having a CAR structure at the data level and either independence, CAR, or ICAR structure at the prior level. For all values of $\rho$ in the data-generating models, the marginal distribution of the observations remains the same for each region due to choosing the same spatially varying coefficient (SVC) vectors. We adopt certain evaluation metrics discussed in Section 6.1 since they collectively capture complementary aspects of the model performance: estimation accuracy, posterior uncertainty, and validity of frequentist properties.
6.1 Evaluation criteria
To evaluate and compare models, we use three complementary metrics: mean squared error (MSE), posterior standard deviation (SD), and coverage probability (CovP). Let $\theta$ denote a generic scalar parameter of interest with true value $\theta^{*}$. Suppose we run $S$ independent simulation replicates. For replicate $s\in\{1,\dots,S\}$ we obtain $M$ posterior draws $\{\theta^{(s)}_{1},\dots,\theta^{(s)}_{M}\}$ (after burn-in and thinning). Suppose we denote the posterior mean and posterior standard deviation in replicate $s$ by
$\hat{\theta}^{(s)}\;=\;\frac{1}{M}\sum_{m=1}^{M}\theta^{(s)}_{m},\qquad\mathrm{SD}^{(s)}(\theta)\;=\;\sqrt{\frac{1}{M-1}\sum_{m=1}^{M}\big(\theta^{(s)}_{m}-\hat{\theta}^{(s)}\big)^{2}}.$
  • Mean Squared Error (MSE): For a parameter $\theta$ with its estimate $\hat{\theta}$, the MSE is $E[(\hat{\theta}-\theta)^{2}]$. In practice, for the simulation study, for each parameter $\theta$, with true value $\theta^{*}$ and posterior mean estimates $\{\hat{\theta}^{(1)},\dots,\hat{\theta}^{(S)}\}$, we compute
    $\mathrm{MSE}(\theta)=\frac{1}{S}\sum_{s=1}^{S}(\hat{\theta}^{(s)}-\theta^{*})^{2}.$
    This metric captures the overall estimation accuracy by combining bias and variance, and is standard in simulation-based model comparisons. Smaller MSE values are preferred, and comparisons of MSE between models indicate which estimator tends to be closer (in squared-error) to the true value on average.
  • Posterior Standard Deviation (SD): In a Bayesian context, the posterior SD reflects the uncertainty in parameter estimates. Comparing average posterior SDs across models provides insight into which model yields more stable and precise inference. For each replicate, we compute the posterior SD, $\mathrm{SD}^{(s)}(\theta)$ from the $M$ posterior draws as above, and then summarize uncertainty across replicates by the average
    $\overline{\mathrm{SD}}(\theta)\;=\;\frac{1}{S}\sum_{s=1}^{S}\mathrm{SD}^{(s)}(\theta).$
    A smaller $\overline{\mathrm{SD}}$ indicates that the model typically yields more precise (less uncertain) posterior estimates. A small posterior SD is desirable only when intervals remain well calibrated; a very small SD combined with poor coverage indicates overconfident inference.
[…]
    $\mathrm{CovP}(\theta)\;=\;\frac{1}{S}\sum_{s=1}^{S}C^{(s)}(\theta).$
    For a nominal $95\%$ interval with $\alpha=0.05$, we expect $\mathrm{CovP}$ close to $0.95$. A larger CovP is preferred because it indicates that posterior credible intervals are well-calibrated in a frequentist sense. Usually, a $\mathrm{CovP}\ll 0.95$ indicates undercoverage (overconfident) and $\mathrm{CovP}\gg 0.95$ indicates overcoverage (possibly inefficient).
In summary, MSE measures estimation accuracy by penalizing squared deviations between parameter estimates and their actual values. Posterior SD summarizes the uncertainty in Bayesian parameter estimates. CovP assesses the calibration of Bayesian credible intervals in a frequentist sense. These criteria provide a balanced framework for evaluating estimation accuracy and uncertainty quantification, central to hierarchical Bayesian modeling.
6.2 Results
@@
For each model, we compare the MSE, posterior SD, and CovP of the parameters $\bm{a},\bm{b},\bm{c}$ and $\rho$ based on the posterior samples drawn using MCMC. For each of the 100 datasets, we obtain MSEs and posterior SDs for each $a_{i}$, $b_{i}$, and $c_{i}$. Here, we report the average of the MSEs for all 34 $a_{i}$s and 100 datasets, for example. For $\rho$, we report the average of the MSEs for all 100 datasets. We also obtain the SDs reported here via a similar averaging. For calculating CovPs, we first calculate the 95% credible interval for the corresponding parameters and check whether the true value falls within that interval. Thus, we obtain a binary value for each dataset and each $i$. Further, we average across all 34 regions and 100 datasets for $a_{i}$s, $b_{i}$s, and $c_{i}$s and report the average. For $\rho$, we only obtain a similar binary value from each dataset and report the average across 100 datasets. We report the results in Table 1. When the true data-generating model has $\rho=0$, we leave the CovPs blank for $\rho$ because the posterior distribution is supported over $(0,1)$ and does not include zero.
@@
When true $\rho=0$, the Independent data layer models yield relatively larger MSEs and smaller coverage probabilities across all parameters, whereas the CAR data layer models clearly improve estimation accuracy. In particular, comparing all the metrics across models for $a_{i}$, $b_{i}$, and $c_{i}$s, we observe that the CAR-ICAR model provides the smallest MSE and the smallest SD with reasonable CovPs. Since $\rho$ is not present in the independent data layer models, we do not report those values for $\rho$. We observe similar MSE and SD values for $\rho$ with different priors with CAR data layer models. In this case, the CAR model with an ICAR prior is more suitable due to the natural spatial pattern of the true SVCs obtained from the Indian rainfall dataset. Here, the posterior variances of the parameters $\rho_{a}$, $\rho_{b}$, and $\rho_{c}$ are high. Hence, even if an ICAR prior is less flexible (due to fixing the spatial autocorrelation parameter to one), it provides a better model fitting. When true $\rho=0.5$, i.e., when the spatial correlation at the data layer is moderate, we see a similar performance comparison. CAR data layer models improve substantially upon Independent models. For $a_{i}$ and $b_{i}$, the CAR-ICAR model again provides the most favorable bias–variance trade-off. For $c_{i}$, the Independent data layer models perform slightly better, showing slightly smaller MSE and posterior SD, although comparable with the CAR-ICAR model. For $\rho$, in terms of MSE, CAR-ICAR models perform well, while for the CAR-Indep model, the average posterior SD is the smallest. When true $\rho=0.9$, i.e., when the spatial correlation at the data layer is strong, Independent models fail to recover the strong spatial correlation. In contrast, the CAR–ICAR and CAR–CAR models consistently outperform them in terms of MSE and SD across all regression parameters, except for the low coverage probability (0.9) for $\rho$ in the case of CAR-ICAR and CAR-CAR models. Based on a global comparison with all data-generating scenarios, the CAR-ICAR model is parsimonious and performs equally or better than the alternatives.
Table 2 reports the average DIC and WAIC values, along with their standard errors, across all simulation replicates. As expected, models with a CAR specification at the data layer consistently achieve lower DIC and WAIC than their Independent counterparts, reflecting the benefit of explicitly modeling spatial correlation among observations. Within the Independent data layer, the choice of prior (Indep, ICAR, or CAR) produces only minor differences, and the CAR data layer models uniformly outperform all such models. Among the latter, the CAR-ICAR model attains the smallest DIC and WAIC in most cases, indicating that combining a CAR copula at the data layer with an ICAR prior on the spatially varying coefficients offers the best trade-off between fit and complexity. The CAR-CAR model exhibits comparable performance, whereas the CAR-Indep model performs slightly worse, particularly in scenarios with stronger spatial dependence. Taken together, the DIC and WAIC results reinforce the earlier findings based on MSE and coverage: the CAR-ICAR model is the most parsimonious and best-performing specification overall.
These findings highlight that the novelty of our approach lies not only in the new combination of gamma regression with CAR copula dependence, but also in its ability to achieve lower estimation uncertainty and better coverage than alternative models that ignore spatial dependence. Thus, the proposed framework offers tangible improvements in statistical performance, which makes it attractive to practitioners beyond the rainfall context.
@@
Table 1: Average MSE, posterior SD, and coverage probability (CovP) of the SVCs and the spatial autocorrelation parameter for independent, CAR, and ICAR priors along with Indep model and CAR model specifications for the data layer. Here, we consider low through high values of the true spatial autocorrelation parameter and report the results averaged across different regions and 100 simulated datasets (for SVCs) and for $\rho$ by averaging across 100 simulated datasets. A smaller MSE and SD value and a higher CovP value are preferred.
Latent
layer
$\longrightarrow$
Indep ICAR CAR
$\rho^{\text{true}}$
Data
layer
$\downarrow$
@@
0 Indep $\bm{a}$ 48.35 4.41 0.96 43.48 4.17 0.96 44.71 4.25 0.97
$\bm{b}$ 1.66 2.79 0.95 1.66 2.78 0.95 1.66 2.79 0.95
$\bm{c}$ 1.16 2.31 0.92 1.07 2.20 0.98 1.11 2.27 0.97
0.5 Indep $\bm{a}$ 47.47 4.40 0.96 42.40 4.15 0.94 43.73 4.24 0.94
$\bm{b}$ 1.71 2.79 0.95 1.71 2.79 0.94 1.71 2.79 0.94
$\bm{c}$ 1.18 2.29 0.86 1.11 2.17 0.91 1.14 2.25 0.91
0.9 Indep $\bm{a}$ 47.88 4.40 0.96 43.65 4.15 0.95 44.69 4.23 0.96
$\bm{b}$ 1.78 2.79 0.89 1.78 2.79 0.89 1.78 2.79 0.90
$\bm{c}$ 1.19 2.29 0.91 1.20 2.13 0.93 1.19 2.22 0.92
0 CAR $\bm{a}$ 48.28 4.41 0.95 43.45 4.16 0.95 44.70 4.24 0.95
$\bm{b}$ 1.67 2.79 0.95 1.65 2.78 0.95 1.66 2.79 0.95
$\bm{c}$ 1.16 2.31 0.95 1.07 2.20 0.95 1.12 2.28 0.95
$\rho$ 5.33 4.06 - 5.11 4.01 - 5.19 4.01 -
0.5 CAR $\bm{a}$ 46.17 4.32 0.94 41.64 4.11 0.95 42.74 4.18 0.95
$\bm{b}$ 1.72 2.79 0.95 1.71 2.79 0.95 1.71 2.79 0.95
$\bm{c}$ 1.22 2.27 0.93 1.15 2.27 0.94 1.18 2.30 0.95
$\rho$ 4.11 4.47 0.94 4.11 4.49 0.96 4.11 4.49 0.96
0.9 CAR $\bm{a}$ 40.09 3.99 0.94 37.30 3.87 0.95 37.97 3.90 0.95
$\bm{b}$ 1.80 2.78 0.94 1.77 2.78 0.94 1.78 2.78 0.94
$\bm{c}$ 1.19 2.24 0.93 1.20 2.40 0.96 1.17 2.34 0.95
$\rho$ 0.55 1.80 0.94 0.55 1.80 0.90 0.55 1.80 0.90
    multiplied by $1000$ for MSE of the parameters $\bm{b},\bm{c}\text{ and }\rho$.
    multiplied by $100$ for SD of the parameters $\bm{b},\bm{c}\text{ and }\rho$.
Table 2: Average DIC and WAIC and their standard error (SE, reported within brackets) for independent, CAR, and ICAR priors, along with Independent and CAR model specification for the data layer. Here, we consider low through high values of the true spatial autocorrelation parameter and report the results averaged across 100 simulated datasets. A smaller DIC and WAIC with a smaller SE value are preferred.
Latent layer $\longrightarrow$ Indep ICAR CAR
$\rho^{\text{true}}$ Data layer $\downarrow$ Criterion
0 Indep DIC 28931.2 (6.73) 28925.9 (6.72) 28926.8 (6.73)
WAIC 28931.9 (6.73) 28926.6 (6.73) 28927.4 (6.73)
0 CAR DIC 28932.3 (6.73) 28927.0 (6.72) 28928.1 (6.72)
WAIC 28933.1 (6.73) 28927.6 (6.72) 28928.7 (6.73)
0.5 Indep DIC 28936.2 (6.94) 28930.5 (6.93) 28931.5 (6.93)
WAIC 28936.9 (6.96) 28932.5 (6.95) 28932.9 (6.94)
0.5 CAR DIC 28844.0 (6.67) 28839.4 (6.66) 28840.1 (6.67)
WAIC 28844.5 (6.69) 28839.9 (6.68) 28840.7 (6.68)
0.9 Indep DIC 28932.6 (9.70) 28924.9 (9.66) 28926.7 (9.67)
WAIC 28935.1 (9.74) 28931.4 (9.73) 28931.6 (9.73)
0.9 CAR DIC 28222.1 (6.87) 28217.6 (6.84) 28218.1 (6.86)
WAIC 28222.9 (6.86) 28218.5 (6.83) 28218.8 (6.85)
@@ 7 Indian rainfall data analysis
We first fit all the models (Indep-Indep, Indep-CAR, Indep-ICAR, CAR-Indep, CAR-CAR, and CAR-ICAR) to the Indian rainfall dataset and compare the models based on specific metrics. Here, unlike the simulation settings, the true values of the parameters are not known, and hence, we cannot compare the models in terms of MSEs and coverage probabilities. Hence, we only compare the average posterior SDs for $a_{i}$s, $b_{i}$s, and $c_{i}$s and the posterior SD of $\rho$. Besides, we compare the DIC and WAIC of the models as described in Section 2.6. We report the results in Table 3. Except for the average posterior SD of $c_{i}$s, all other metrics indicate similar or superior performance of the CAR-ICAR model. Regarding DIC and WAIC, the ICAR prior provides a better fitting and prediction performance than the CAR prior, followed by the independent priors; this result indicates the suitability of allowing SVCs in a Gaussian graphical model framework. The observed differences in DIC (exceeding 5) [spiegelhalter2002bayesian] and WAIC (exceeding 11) are regarded as practically meaningful [Gelman2014understanding], thereby further supporting the preference for the CAR-ICAR specification. In contrast, the CAR-CAR model provides a similar fit but with higher uncertainty for a few parameters, suggesting possible overparameterization.
@@
Model $\frac{1}{n}\sum_{i}\mathrm{sd}(a_{i})$ $\frac{1}{n}\sum_{i}\mathrm{sd}(b_{i})$ $\frac{1}{n}\sum_{i}\mathrm{sd}(c_{i})$ $\mathrm{sd}(\rho)$ DIC WAIC
Indep-Indep 4.447 2.722 1.997 - 29042 29049
Indep-ICAR 4.207 2.719 1.865 - 29033 29044
Indep-CAR 4.301 2.701 1.939 - 29037 29045
CAR-Indep 4.105 2.753 2.077 1.225 28141 28155
CAR-ICAR 3.887 2.786 2.293 1.141 28133 28144
CAR-CAR 3.983 2.844 2.210 1.198 28136 28148
    multiplied by $100$ for the parameters $\bm{b},\bm{c}\text{ and }\rho$.
To assess absolute model fit in addition to relative model selection criteria, we perform posterior predictive checks [gelman1996posterior]. For each posterior draw of the parameters, we simulate replicated datasets from the posterior predictive distribution and compute a set of discrepancy statistics: overall mean, standard deviation (SD), minimum, and maximum, average of site-specific means, standard deviations, minimums and maximums, and average neighbor correlation across years. Posterior predictive $p$-values [meng1994posterior] are calculated as the proportion of replicated statistics more extreme than the observed. These posterior predictive $p$-values are reported in Table 4. Across all models, the $p$-values for the mean, SDs, minimum, maximum, site-level summaries, and neighbor correlation range between 0.14 and 0.95, indicating no evidence of lack of fit for these global summaries. The CAR-ICAR model remains the best-supported specification, combining superior relative fit (WAIC/DIC) with adequate posterior predictive performance for key distributional and spatial summaries.
Table 4: Posterior predictive $p$-value analysis for different models using different types of test statistics. A $p$-value more than $0.95$ or less than $0.05$ indicates that the model misfits the data. In an ideal case, the $p$-value is approximately $0.5$. In the bottom, we reported the coverage probability of the data, which measures the proportion of how many times the observed data falls within the $95\%$ posterior predictive data samples.
Test statistic $\downarrow$
Indep-
Indep
Indep-
ICAR
Indep-
CAR
CAR-
Indep
CAR-
ICAR
CAR-
CAR
Mean 0.58 0.56 0.58 0.52 0.66 0.61
SD 0.55 0.55 0.55 0.80 0.89 0.86
Minimum 0.67 0.65 0.65 0.85 0.85 0.85
Maximum 0.36 0.37 0.37 0.53 0.61 0.59
Site wise mean 0.53 0.51 0.53 0.50 0.64 0.59
Site wise SD 0.67 0.68 0.68 0.91 0.95 0.93
Site wise minimum 0.29 0.28 0.29 0.14 0.17 0.17
Site wise maximum 0.43 0.43 0.43 0.71 0.83 0.78
Avg neighbor correlation - - - 0.64 0.73 0.66
Coverage Probability 0.96 0.96 0.96 0.95 0.95 0.96
@@
We further consider the CAR-ICAR model only for drawing inferences and report the results based on it. To assess the quality of MCMC chains, we utilize Geweke statistics for the convergence of the post-burn-in samples and the effective sample size (ESS) for mixing and sample autocorrelation. We report the average of the Geweke statistics, and we use the batch-mean methodology [brooks2011handbook] to obtain the ESS values for SVCs. A high ESS value is preferred. The Geweke statistic is a $z$-score; hence, its absolute value being lower than 1.96 indicates the convergence of the MCMC chains. We report the results in Table 5. While the ESS values are relatively higher for all the model hyperparameters, they are smaller for the model parameter $\rho$. Overall, the ESS values are reasonable for all parameters and hyperparameters. Further, the Geweke statistics for all parameters and hyperparameters are close to zero and not outside the interval $\pm 1.96$, indicating convergence of the MCMC chains after the burn-in periods. Altogether, both metrics suggest that the MCMC samples are of reasonably good quality for drawing statistical inferences. Convergence of the MCMC algorithm is assessed using univariate trace plots of representative parameters across all models. The chains display stable mixing and no discernible trends. The complete set of trace plots and the ESS table of the parameters, considering all other models, are reported in Section 3 of the Supplementary Material. For the Indian monsoon data, posterior predictive checks similarly suggest that the CAR-Copula models provide a satisfactory fit. The $p$-values remain in reasonable ranges and do not indicate a systematic lack of fit. Section 4 of Supplementary Material includes further posterior predictive plots and diagnostics supporting these findings.
@@
Table 5: Average ESS calculated using batch-means and Geweke statistics of different parameters of the CAR-ICAR model
@@
We report the posterior means and SDs of $a_{i}$s, $b_{i}$s, and $c_{i}$s in Figure 3. Besides, the posterior mean and SD of $\rho$ are 0.9309 and 0.0114, respectively. The spatial patterns of the coefficients are similar to the ones observed in Figure 2, which indicates that the CAR-ICAR model can capture the spatial patterns of the coefficients correctly. The posterior means of $a_{i}$s, i.e., the shape parameters of the gamma model, are higher in the middle, eastern, and southern parts of India. At the same time, they are low in the northwestern parts of India and also near the northeastern meteorological subdivisions. Several eastern and western Indian subdivisions showcase higher posterior SD than others. The posterior means of $b_{i}$s, the intercept-related parameters of the gamma model, are higher in the northwestern and northeastern meteorological subdivisions. At the same time, they are lowest in the middle parts of the country. Finally, a positive $c_{i}$ indicates a negative trend follows from (11). We observe positive estimates of $c_{i}$s throughout the country, indicating a negative trend in the amount of monsoon total rainfall. However, due to high posterior SDs, most of these estimates are insignificant except for three meteorological subdivisions: 1. Arunachal Pradesh, 2. Assam and Meghalaya, and 3. Nagaland, Manipur, Mizoram, and Tripura. A significantly negative trend for these Himalayan regions indicates a decline in mean total monsoon rainfall, and thus requires further environmental assessment.